Molecular Structure Determination via SnB Russ Miller

Department of Computer Science, University at Buffalo Department of Molecular Biophysics, Hauptman-Woodward Medical

Research Institute miller@cs.buffalo.edu

SnB is a direct-methods computer program based on Shake-and-Bake, an ab initio, dual-space, multiple-solution procedure capable of producing automated solutions for molecular crystal structures, provided that atomicresolution diffraction data are available. This iterative process incorporates reciprocal-space optimization (phase refinement) with filtering in real-space to impose constraints. Phase refinement is performed by either (i) a parameter shift routine, which makes several passes through the phase set and adjusts the phases, in sequence, so as to reduce the value of Hauptman’s Minimal Function, or (ii) the tangent formula, in which case the Minimal Function acts passively as a figure-of-merit. A simulated-annealing approach incorporating aspects of Shake-and-Bake also shows promise if the parameters are tuned well.

SnB is supported on a variety of platforms, including Unix workstations, the Thinking Machines Corporation CM-5, the Cray C90, and the new massively parallel Cray T3D. Information about SnB, including availability, can be obtained from the WWW home page located at http://www.hwi.buffalo.edu/SnB/.

The public version of SnB Vers. 1.0 has been used to determine a number of difficult unknown structures within the past year. These structures had resisted traditional approaches, some for as many as 7 years. These include a 144 non-H atom P1 structure that was solved at the Scripps Research Institute, a 188 non-H atom P1 structure that was solved at Molecular Structure Corporation, a 300 non-H atom structure that was solved in a lab at UCLA, and a 400 non-H atom structure that was quite recently solved in a lab at the University of Toronto. Several smaller unknown structures have also been solved recently.

This is joint work with Chuck Weeks, Herb Hauptman, and Steve Gallo. Research support from NIH and NSF.